OpenStudy (anonymous):

solve the inequality |x-8|>22

OpenStudy (saifoo.khan):

|x-8| > 22 |x-8| < -22

OpenStudy (saifoo.khan):

Joathday, am i correct?

OpenStudy (anonymous):

noo,,

OpenStudy (saifoo.khan):

Oh, where did i go wrong?

OpenStudy (anonymous):

when u split the equation u take away the absolute value. other than that its fine just not finished

OpenStudy (anonymous):

Question: 18 of 20: Solve the inequality: Select one of the options below as your answer: A. x > -30 or x < -14 B. x > 30 or x < -14 C. x > 30 or x < 14 D. x > 30 and x < -14

OpenStudy (saifoo.khan):

Ok, so its like this: x-8 > 22 x-8 < -22

OpenStudy (anonymous):

x>30 or x<-14 and yes

OpenStudy (saifoo.khan):

its B

OpenStudy (saifoo.khan):

right?

OpenStudy (saifoo.khan):

Yeas, it is! :D

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

Which number line represents the solution set of the equation Select one of the options below as your answer |x+6|=2 where would the point go

OpenStudy (saifoo.khan):

x = -4, i guess.

OpenStudy (anonymous):

x=-8,-4

OpenStudy (anonymous):

Use the quadratic formula to solve the quadratic equation: 2x2 − 3x − 7 = 0

OpenStudy (saifoo.khan):

Alright, got that! ^.^

OpenStudy (anonymous):

\[(3\pm \sqrt{65})/4\]

OpenStudy (anonymous):

i dont have a calculator with me so thats the best im gonna do

OpenStudy (anonymous):

thanks is it + eqals to or just +

OpenStudy (anonymous):

its plus or minus which means there is possibly two answers. in 1 equation u have to add them and in another u have to subtract. i would check it for u if i had a cal but sadly i wont have it for another hour as my brother uses it